# MDReID: Modality-Decoupled Learning for Any-to-Any Multi-Modal Object Re-Identification

### v1

@@ 1 Introduction
Object Re-Identification (ReID) focuses on identifying and retrieving specific objects across non-overlapping camera views. Conventional object re-identification (ReID) methods [1, 2, 3, 4, 5, 6, 7] primarily rely on RGB images. However, RGB-based approaches often face significant limitations under challenging environmental conditions, such as poor illumination, shadows, and low image resolution. These adverse conditions frequently lead to the extraction of misleading features, thereby diminishing discriminative capability. To overcome these limitations, multi-modal ReID [8, 9, 10, 11, 12, 13], which integrates complementary information from multiple spectrums, has emerged as a promising approach. By leveraging the inherent strengths and complementary features of different modalities, multi-modal ReID significantly enhances feature representations, enabling more robust and accurate identification in complex real-world scenarios.
@@
The introduction of multi-modal data has significantly enhanced the performance of ReID in challenging scenarios, thereby stimulating extensive research efforts in this field [8, 9, 10, 11, 12]. EDITOR [11] selects diverse tokens from Vision Transformer (ViT) to mitigate the impact of irrelevant backgrounds and narrow the gap between modalities, achieving strong detection performance. TOP-ReID [12] incorporates a cyclic token permutation module to aggregate multi-spectral features and a complementary reconstruction module to minimize the distribution gap across different image spectra. These innovations enable TOP-ReID to achieve robust performance under both modality-complete and modality-missing scenarios. Although these methods effectively leverage the characteristics of multi-modal data, they are developed based on a fundamental assumption that all modalities within the dataset are strictly aligned (as illustrated in Figure 1 (a) Modality-matched). Ideally, every camera would simultaneously provide RGB, NIR, and TIR modalities. However, due to practical constraints such as device differences and deployment diversity, achieving fully matched retrieval conditions is challenging in real-world scenarios. Therefore, it is crucial to develop a flexible image-level any-to-any ReID framework capable of effectively operating under both modality-matched and modality-mismatched scenarios (as illustrated in Figure 1 (a)).
To address this limitation, we propose MDReID, a flexible any-to-any ReID framework designed to support retrieval tasks involving arbitrary combinations of query and gallery modalities. The core challenge of this problem lies in learning effective representations that remain robust under both modality-matched and modality-mismatched conditions. A straightforward solution, as adopted by TOP-ReID [12], attempts to predict the missing modality representation from the available one, thereby enabling modality-aligned retrieval. However, as highlighted by RLE [14], predicting cross-spectrum features solely from visual inputs constitutes an ill-posed problem. The modality-specific characteristics that are inherently unpredictable often lead to suboptimal learning. To this end, MDReID introduces a Modality-Decoupled Learning (MDL), which introduces a modality-shared and modality-specific token into the ViT architecture, leveraging multi-layer attention within the transformer to extract shared and specific features from multi-modal inputs. MDL explicitly disentangles the modality representations into two complementary components: a modality-shared component that captures the predictable cross-modal representation and a modality-specific component that preserves the unpredictable modality characteristics. This decoupling enables the model to better generalize across diverse modality combinations while maintaining the modality-specific cues.
@@ 2 Related Work
Compared to the general RGB-to-RGB object re-identification [15, 16, 17, 18, 19, 20, 21], multi-modal object re-identification (ReID) has demonstrated significant practical value, achieving widespread applications in cross-scenario systems [22, 23] (e.g., security surveillance, intelligent transportation, etc.). Current methods [24, 9, 25] primarily enhance re-identification accuracy by effectively leveraging the complementarity and integration of multi-modal features. For example, Zheng et al. [26] proposed the cross-directional consistency network (CCNet), which overcomes discrepancies in both modality and sample aspects, thereby achieving more effective multi-modal data collaboration. To address the modal-missing problem, Zheng et al. [27] proposed the DENet model, which incorporates a feature transformation module to recover information from missing modalities and a dynamic enhancement module to improve multi-modality representation. Li et al. [8] proposed HAMNet, which automatically fuses different spectral features using a specially designed heterogeneous score coherence loss. He et al. [28] proposed the GPFNet model, which utilizes graph learning to fuse multi-modal features.
Notably, models based on Vision Transformer (ViT) [29, 30, 10, 31, 32, 33] for object re-identification have achieved excellent results in recent years and gained widespread attention. Pan et al. [34] proposed the H-ViT model, which alleviates challenges caused by heterogeneous multi-modalities and reduces feature deviations arising from modal variations. Additionally, they introduced a progressively hybrid transformer (PHT [35]) that effectively fuses multi-modal complementary information through random hybrid augmentation and a feature hybrid mechanism. Crawford et al. [36] addressed the issue of modality laziness in multi-modal fusion by proposing the UniCat model based on a ViT architecture. By selecting diverse tokens from ViT to mitigate the impact of irrelevant backgrounds and narrow the gap between modalities, Zhang et al. [11] developed the EDITOR model. Wang et al. [12] designed the Top-ReID model, which achieves state-of-the-art accuracy by incorporating a cyclic token permutation module to aggregate multi-spectral features and a complementary reconstruction module to minimize the distribution gap across different image spectra.
@@ 4.1 Datasets and Evaluation Protocols
We evaluate the proposed MDReID framework on three publicly available datasets spanning both person (RGBNT201 [24]) and vehicle re-identification (RGBNT100 [8] and MSVR310 [26]) tasks. For evaluating modality mismatch scenarios, we focus on representative cross-modal configurations, including RT-to-NT, RT-to-N, R-to-N, and R-to-NT. Here, NIR is the second most widely used modality in surveillance systems after RGB, and RGB-to-NIR research continues to attract significant interest. Therefore, we evaluate our method in four R-to-N-based application scenarios. Performance is measured using mean Average Precision (mAP) and Cumulative Matching Characteristics (CMC) at ranks 1, 5, and 10 (R-1, R-5, R-10). We also report key complexity indicators, including the number of trainable parameters and floating point operations (FLOPs), to quantify the efficiency of the model.
RGBNT201 [24] is a multi-modality person re-identification dataset designed to overcome the limitations of single-modal imaging in challenging surveillance scenarios. It was captured on a university campus using four non-overlapping camera views with a synchronized triple-camera system that simultaneously records RGB, near-infrared (NIR), and thermal-infrared (TIR) images. The dataset contains 201 identities, each represented by at least 20 non-adjacent image triplets, resulting in 4,787 images per modality. The collection covers a range of real-world conditions, including severe illumination changes, occlusion, viewpoint variations, and background clutter. The dataset is split into 141 identities for training, 30 for validation, and 30 for testing, with the entire test set serving as the gallery and 10 records per identity sampled as probes. RGBNT201 thus provides a robust benchmark for evaluating multi-modal fusion strategies and addressing missing-modality issues in person re-identification.
@@
MSVR310 [26] is a high-quality multi-spectral vehicle re-identification benchmark captured under diverse, challenging conditions. It contains 2,087 triplet samples (6,261 images in total) from 310 vehicles. Each sample includes spatially aligned images from three modalities: RGB, NIR, and TIR. RGB images are captured by a 360 D866 camera during the day and by a Mi8 mobile phone at night, NIR images are obtained with the 360 D866 in near-infrared mode, and TIR images are recorded with a FLIR SC620 camera at 640×480 resolution. The number of samples per vehicle ranges from 2 to 20, and time labels are provided for cross-time matching. The dataset is divided into 1,032 samples from 155 vehicles for training and 1,055 samples from 155 vehicles for testing/gallery, with 591 samples (from 52 vehicles) used as queries. MSVR310 provides a comprehensive platform for addressing intra-class appearance variations and modality differences to support robust vehicle re-identification.
@@ 4.2 Implementation Details
All experiments are implemented in PyTorch and conducted on a single NVIDIA RTX 4090 GPU with CUDA 12.5 and Python 3.8. We adopt the CLIP-Base [37] visual encoder as the backbone. Input images are resized to $256\times 128$ for RGBNT201 and $128\times 256$ for RGBNT100 and MSVR310. We employ random horizontal flipping, cropping, and erasing [38] for data augmentation. The model is optimized using Adam with a batch size of 64. The base learning rate is initialized at $3.5\times 10^{-4}$, while the visual encoder is fine-tuned with a reduced rate of $5\times 10^{-6}$. Training is performed for 50 epochs.
@@
MUDeep [39] 23.8 19.7 33.1 44.3 DMML [40] 58.5 82.0 19.1 31.1
HACNN [41] 21.3 19.0 34.1 42.8 BoT [42] 78.0 95.1 23.5 38.4
MLFN [43] 26.1 24.2 35.9 44.1 Circle Loss [44] 59.4 81.7 22.7 34.2
PCB [45] 32.8 28.1 37.4 46.9 HRCN [46] 67.1 91.8 23.4 44.2
OSNet [47] 25.4 22.3 35.1 44.7 TransReID [1] 75.6 92.9 18.4 29.6
CAL [48] 27.6 24.3 36.5 45.7 AGW [16] 73.1 92.7 28.9 46.9
@@
HAMNet [8] 27.7 26.3 41.5 51.7 GAFNet [25] 74.4 93.4 - -
PFNet [24] 38.5 38.9 52.0 58.4 GraFT [49] 76.6 94.3 - -
IEEE [9] 47.5 44.4 57.1 63.6 GPFNet [28] 75.0 94.5 - -
DENet [27] 42.4 42.2 55.3 64.5 PHT [35] 79.9 92.7 - -
UniCat [36] 57.0 55.7 - - UniCat [36] 79.4 96.2 - -
HTT [10] 71.1 73.4 83.1 87.3 CCNet [26] 77.2 96.3 36.4 55.2
EDITOR [11] 66.5 68.3 81.1 88.2 EDITOR [11] 82.1 96.4 39.0 49.3
RSCNet [31] 68.2 72.5 - - RSCNet [31] 82.3 96.6 39.5 49.6
TOP-ReID [12] 72.3 76.6 84.7 89.4 TOP-ReID [12] 81.2 96.4 35.9 44.6
@@ 4.3 Comparison with State-of-the-Art Methods
Multi-spectral Object ReID (RNT-to-RNT) We compare our MDReID against a range of state-of-the-art single-spectral and multi-spectral approaches on RGBNT201, RGBNT100, and MSVR310. For person ReID on RGBNT201, as shown in Table 1, the TOP-ReID [12] achieves 72.3% mAP, 76.6% R-1, 84.7% R-5, and 89.4% R-10. In comparison, our MDReID outperforms these results by 9.8%, 8.6%, 5.6%, and 3.2%, respectively, highlighting the effectiveness of our proposed MDReID. For vehicle re-ID, MDReID achieves the highest mAP on RGBNT100 (85.3%) and surpasses the next-best approach on MSVR310 by a significant margin of 11.5% in mAP and 13.7% in R-1. These results clearly demonstrate the superior discriminative power and robustness of our modality-decoupled design in multi-spectral re-ID scenarios where all modalities are available.
Evaluation on Missing-spectral Scenarios. In practical applications, sensor limitations or environmental factors often lead to incomplete modality inputs. Following the setting of TOP-ReID [12], we evaluate the robustness of our method under all six possible missing-modality configurations across the RGB, NIR, and TIR channels using the RGBNT201 dataset. As shown in Table 2, our MDReID consistently outperforms the TOP-ReID, achieving an average improvement of 10.0% in mAP and 9.4% in Rank-1 accuracy across all missing-modality scenarios. These results demonstrate that our modality-decoupled framework remains highly effective even when partial modality information is absent, highlighting its strong generalization ability in incomplete multispectral settings.
@@
Single
MUDeep [39] 19.2 16.4 20.0 17.2 18.4 14.2 13.7 11.8 11.5 6.5 12.7 8.5 15.9 12.9
HACNN [41] 12.5 11.1 20.5 19.4 16.7 13.3 9.2 6.2 6.3 2.2 14.8 12.0 13.3 10.7
MLFN [43] 20.2 18.9 21.1 19.7 17.6 11.1 13.2 12.1 8.3 3.5 13.1 9.1 15.6 12.4
PCB [45] 23.6 24.2 24.4 25.1 19.9 14.7 20.6 23.6 11.0 6.8 18.6 14.4 19.7 18.1
OSNet [47] 19.8 17.3 21.0 19.0 18.7 14.6 12.3 10.9 9.4 5.4 13.0 10.2 15.7 12.9
Multi
PFNet [24] - - 31.9 29.8 25.5 25.8 - - - - 26.4 23.4 - -
DENet [27] - - 35.4 36.8 33.0 35.4 - - - - 32.4 29.2 - -
TOP-ReID [12] 54.4 57.5 64.3 67.6 51.9 54.5 35.3 35.4 26.2 26.0 34.1 31.7 44.4 45.4
@@
Evaluation on Modality-mismatched Scenarios. To evaluate the generalizability of our approach under realistic deployment conditions, we benchmark MDReID across four representative modality-mismatched scenarios on three public datasets. We compare against two strong baselines, TOP-ReID [12] and EDITOR [11], by reproducing their results using official open-source implementations. As shown in Table 3, EDITOR is designed exclusively for modality-matched settings, which exhibits limited effectiveness when modalities differ across query and gallery. Although TOP-ReID attempts to address missing modalities via reconstruction, its performance remains constrained due to the ill-posed nature of the reconstruction problem. In contrast, our MDReID framework consistently improves the average mAP by 3.4%, 11.8%, and 10.9% on the RGBNT201, RGBNT100, and MSVR310 datasets, respectively. In addition, we train four expert models on the RGBNT201 dataset for four specific scenarios. Results show that our method outperforms these experts by 9.2% on average in mAP and by 8.9% in R‑1. Importantly, our single‑model approach adapts to all four scenarios and offers greater flexibility. In summary, these results highlight the robustness and adaptability of MDReID in handling modality-mismatched ReID, affirming its effectiveness capable of supporting arbitrary modality combinations in both query and gallery inputs.
@@
RGBNT201 EDITOR [11] 27.3 27.9 2.80 0.0 4.0 2.3 4.3 4.1 8.5 7.5
TOP-ReID [12] 43.0 44.6 14.4 13.3 15.4 14.0 11.9 8.7 18.2 18.0
@@
RGBNT100 EDITOR [11] 42.1 59.9 2.6 0.8 2.8 1.5 2.7 1.5 11.9 15.5
TOP-ReID [12] 59.0 81.8 21.9 28.5 26.2 34.0 21.7 25.4 26.8 36.1
@@
MSVR310 EDITOR [11] 6.4 11.0 2.2 2.5 1.6 0.2 1.7 1.5 2.5 3.4
TOP-ReID [12] 18.4 30.5 12.9 19.5 13.7 21.3 12.6 18.8 11.2 17.8

### v2

@@ 1 Introduction
Object Re-Identification (ReID) focuses on identifying and retrieving specific objects across non-overlapping camera views. Conventional object re-identification (ReID) methods He et al. (2021); Li et al. (2023b); Zhang et al. (2023); Li et al. (2023a); Wang et al. (2024a); Tan et al. (2022, 2024a) primarily rely on RGB images. However, RGB-based approaches often face significant limitations under challenging environmental conditions, such as poor illumination, shadows, and low image resolution. These adverse conditions frequently lead to the extraction of misleading features, thereby diminishing discriminative capability. To overcome these limitations, multi-modal ReID Li et al. (2020); Wang et al. (2022, 2024d); Zhang et al. (2024); Wang et al. (2024c); Feng et al. , which integrates complementary information from multiple spectrums, has emerged as a promising approach. By leveraging the inherent strengths and complementary features of different modalities, multi-modal ReID significantly enhances feature representations, enabling more robust and accurate identification in complex real-world scenarios.
@@
The introduction of multi-modal data has significantly enhanced the performance of ReID in challenging scenarios, thereby stimulating extensive research efforts in this field Li et al. (2020); Wang et al. (2022, 2024d); Zhang et al. (2024); Wang et al. (2024c). EDITOR Zhang et al. (2024) selects diverse tokens from Vision Transformer (ViT) to mitigate the impact of irrelevant backgrounds and narrow the gap between modalities, achieving strong detection performance. TOP-ReID Wang et al. (2024c) incorporates a cyclic token permutation module to aggregate multi-spectral features and a complementary reconstruction module to minimize the distribution gap across different image spectra. These innovations enable TOP-ReID to achieve robust performance under both modality-complete and modality-missing scenarios. Although these methods effectively leverage the characteristics of multi-modal data, they are developed based on a fundamental assumption that all modalities within the dataset are strictly aligned (as illustrated in Figure 1 (a) Modality-matched). Ideally, every camera would simultaneously provide RGB, NIR, and TIR modalities. However, due to practical constraints such as device differences and deployment diversity, achieving fully matched retrieval conditions is challenging in real-world scenarios. Therefore, it is crucial to develop a flexible image-level any-to-any ReID framework capable of effectively operating under both modality-matched and modality-mismatched scenarios (as illustrated in Figure 1 (a)).
To address this limitation, we propose MDReID, a flexible any-to-any ReID framework designed to support retrieval tasks involving arbitrary combinations of query and gallery modalities. The core challenge of this problem lies in learning effective representations that remain robust under both modality-matched and modality-mismatched conditions. A straightforward solution, as adopted by TOP-ReID Wang et al. (2024c), attempts to predict the missing modality representation from the available one, thereby enabling modality-aligned retrieval. However, as highlighted by RLE Tan et al. (2024c), predicting cross-spectrum features solely from visual inputs constitutes an ill-posed problem. The modality-specific characteristics that are inherently unpredictable often lead to suboptimal learning. To this end, MDReID introduces a Modality-Decoupled Learning (MDL), which introduces a modality-shared and modality-specific token into the ViT architecture, leveraging multi-layer attention within the transformer to extract shared and specific features from multi-modal inputs. MDL explicitly disentangles the modality representations into two complementary components: a modality-shared component that captures the predictable cross-modal representation and a modality-specific component that preserves the unpredictable modality characteristics. This decoupling enables the model to better generalize across diverse modality combinations while maintaining the modality-specific cues.
@@ 2 Related Work
Compared to the general RGB-to-RGB object re-identification Li et al. (2021); Ye et al. (2022); Wang et al. (2025a); Tan et al. (2024b); Jia et al. (2025); Yin et al. (2025); Liu et al. (2024), multi-modal object re-identification (ReID) has demonstrated significant practical value, achieving widespread applications in cross-scenario systems Su et al. (2025); Zhang and Wang (2023a); Wang et al. (2025b, c) (e.g., security surveillance, intelligent transportation, etc.). Current methods Zheng et al. (2021); Wang et al. (2022); Guo et al. (2022) primarily enhance re-identification accuracy by effectively leveraging the complementarity and integration of multi-modal features. For example, Zheng et al. Zheng et al. (2023b) proposed the cross-directional consistency network (CCNet), which overcomes discrepancies in both modality and sample aspects, thereby achieving more effective multi-modal data collaboration. To address the modal-missing problem, Zheng et al. Zheng et al. (2023a) proposed the DENet model, which incorporates a feature transformation module to recover information from missing modalities and a dynamic enhancement module to improve multi-modality representation. Li et al. Li et al. (2020) proposed HAMNet, which automatically fuses different spectral features using a specially designed heterogeneous score coherence loss. He et al. He et al. (2023) proposed the GPFNet model, which utilizes graph learning to fuse multi-modal features.
Notably, models based on Vision Transformer (ViT) Ye et al. (2024); Wang et al. (2024b, d); Yu et al. (2024); Li et al. (2024); Zhu et al. (2023) for object re-identification have achieved excellent results in recent years and gained widespread attention. Pan et al. Pan et al. (2022) proposed the H-ViT model, which alleviates challenges caused by heterogeneous multi-modalities and reduces feature deviations arising from modal variations. Additionally, they introduced a progressively hybrid transformer (PHT Pan et al. (2023)) that effectively fuses multi-modal complementary information through random hybrid augmentation and a feature hybrid mechanism. Crawford et al. Crawford et al. (2023) addressed the issue of modality laziness in multi-modal fusion by proposing the UniCat model based on a ViT architecture. By selecting diverse tokens from ViT to mitigate the impact of irrelevant backgrounds and narrow the gap between modalities, Zhang et al. Zhang et al. (2024) developed the EDITOR model. Wang et al. Wang et al. (2024c) designed the Top-ReID model, which achieves state-of-the-art accuracy by incorporating a cyclic token permutation module to aggregate multi-spectral features and a complementary reconstruction module to minimize the distribution gap across different image spectra.
@@ 4.1 Datasets and Evaluation Protocols
We evaluate the proposed MDReID framework on three publicly available datasets spanning both person (RGBNT201 Zheng et al. (2021)) and vehicle re-identification (RGBNT100 Li et al. (2020) and MSVR310 Zheng et al. (2023b)) tasks. For evaluating modality mismatch scenarios, we focus on representative cross-modal configurations, including RT-to-NT, RT-to-N, R-to-N, and R-to-NT. Here, NIR is the second most widely used modality in surveillance systems after RGB, and RGB-to-NIR research continues to attract significant interest. Therefore, we evaluate our method in four R-to-N-based application scenarios. Performance is measured using mean Average Precision (mAP) and Cumulative Matching Characteristics (CMC) at ranks 1, 5, and 10 (R-1, R-5, R-10). We also report key complexity indicators, including the number of trainable parameters and floating point operations (FLOPs), to quantify the efficiency of the model.
RGBNT201 Zheng et al. (2021) is a multi-modality person re-identification dataset designed to overcome the limitations of single-modal imaging in challenging surveillance scenarios. It was captured on a university campus using four non-overlapping camera views with a synchronized triple-camera system that simultaneously records RGB, near-infrared (NIR), and thermal-infrared (TIR) images. The dataset contains 201 identities, each represented by at least 20 non-adjacent image triplets, resulting in 4,787 images per modality. The collection covers a range of real-world conditions, including severe illumination changes, occlusion, viewpoint variations, and background clutter. The dataset is split into 141 identities for training, 30 for validation, and 30 for testing, with the entire test set serving as the gallery and 10 records per identity sampled as probes. RGBNT201 thus provides a robust benchmark for evaluating multi-modal fusion strategies and addressing missing-modality issues in person re-identification.
@@
MSVR310 Zheng et al. (2023b) is a high-quality multi-spectral vehicle re-identification benchmark captured under diverse, challenging conditions. It contains 2,087 triplet samples (6,261 images in total) from 310 vehicles. Each sample includes spatially aligned images from three modalities: RGB, NIR, and TIR. RGB images are captured by a 360 D866 camera during the day and by a Mi8 mobile phone at night, NIR images are obtained with the 360 D866 in near-infrared mode, and TIR images are recorded with a FLIR SC620 camera at 640×480 resolution. The number of samples per vehicle ranges from 2 to 20, and time labels are provided for cross-time matching. The dataset is divided into 1,032 samples from 155 vehicles for training and 1,055 samples from 155 vehicles for testing/gallery, with 591 samples (from 52 vehicles) used as queries. MSVR310 provides a comprehensive platform for addressing intra-class appearance variations and modality differences to support robust vehicle re-identification.
@@ 4.2 Implementation Details
All experiments are implemented in PyTorch and conducted on a single NVIDIA RTX 4090 GPU with CUDA 12.5 and Python 3.8. We adopt the CLIP-Base Radford et al. (2021) visual encoder as the backbone. Input images are resized to $256\times 128$ for RGBNT201 and $128\times 256$ for RGBNT100 and MSVR310. We employ random horizontal flipping, cropping, and erasing Zhong et al. (2020) for data augmentation. The model is optimized using Adam with a batch size of 64. The base learning rate is initialized at $3.5\times 10^{-4}$, while the visual encoder is fine-tuned with a reduced rate of $5\times 10^{-6}$. Training is performed for 50 epochs.
@@
MUDeep Qian et al. (2017) 23.8 19.7 33.1 44.3 DMML Chen et al. (2019) 58.5 82.0 19.1 31.1
HACNN Li et al. (2018) 21.3 19.0 34.1 42.8 BoT Luo et al. (2019) 78.0 95.1 23.5 38.4
MLFN Chang et al. (2018) 26.1 24.2 35.9 44.1 Circle Loss Sun et al. (2020) 59.4 81.7 22.7 34.2
PCB Sun et al. (2018) 32.8 28.1 37.4 46.9 HRCN Zhao et al. (2021) 67.1 91.8 23.4 44.2
OSNet Zhou et al. (2019) 25.4 22.3 35.1 44.7 TransReID He et al. (2021) 75.6 92.9 18.4 29.6
CAL Rao et al. (2021) 27.6 24.3 36.5 45.7 AGW Ye et al. (2022) 73.1 92.7 28.9 46.9
@@
HAMNet Li et al. (2020) 27.7 26.3 41.5 51.7 GAFNet Guo et al. (2022) 74.4 93.4 - -
PFNet Zheng et al. (2021) 38.5 38.9 52.0 58.4 GraFT Yin et al. (2023) 76.6 94.3 - -
IEEE Wang et al. (2022) 47.5 44.4 57.1 63.6 GPFNet He et al. (2023) 75.0 94.5 - -
DENet Zheng et al. (2023a) 42.4 42.2 55.3 64.5 PHT Pan et al. (2023) 79.9 92.7 - -
UniCat Crawford et al. (2023) 57.0 55.7 - - UniCat Crawford et al. (2023) 79.4 96.2 - -
HTT Wang et al. (2024d) 71.1 73.4 83.1 87.3 CCNet Zheng et al. (2023b) 77.2 96.3 36.4 55.2
EDITOR Zhang et al. (2024) 66.5 68.3 81.1 88.2 EDITOR Zhang et al. (2024) 82.1 96.4 39.0 49.3
RSCNet Yu et al. (2024) 68.2 72.5 - - RSCNet Yu et al. (2024) 82.3 96.6 39.5 49.6
TOP-ReID Wang et al. (2024c) 72.3 76.6 84.7 89.4 TOP-ReID Wang et al. (2024c) 81.2 96.4 35.9 44.6
@@ 4.3 Comparison with State-of-the-Art Methods
Multi-spectral Object ReID (RNT-to-RNT) We compare our MDReID against a range of state-of-the-art single-spectral and multi-spectral approaches on RGBNT201, RGBNT100, and MSVR310. For person ReID on RGBNT201, as shown in Table 1, the TOP-ReID Wang et al. (2024c) achieves 72.3% mAP, 76.6% R-1, 84.7% R-5, and 89.4% R-10. In comparison, our MDReID outperforms these results by 9.8%, 8.6%, 5.6%, and 3.2%, respectively, highlighting the effectiveness of our proposed MDReID. For vehicle re-ID, MDReID achieves the highest mAP on RGBNT100 (85.3%) and surpasses the next-best approach on MSVR310 by a significant margin of 11.5% in mAP and 13.7% in R-1. These results clearly demonstrate the superior discriminative power and robustness of our modality-decoupled design in multi-spectral re-ID scenarios where all modalities are available.
Evaluation on Missing-spectral Scenarios. In practical applications, sensor limitations or environmental factors often lead to incomplete modality inputs. Following the setting of TOP-ReID Wang et al. (2024c), we evaluate the robustness of our method under all six possible missing-modality configurations across the RGB, NIR, and TIR channels using the RGBNT201 dataset. As shown in Table 2, our MDReID consistently outperforms the TOP-ReID, achieving an average improvement of 10.0% in mAP and 9.4% in Rank-1 accuracy across all missing-modality scenarios. These results demonstrate that our modality-decoupled framework remains highly effective even when partial modality information is absent, highlighting its strong generalization ability in incomplete multispectral settings.
@@
Single MUDeep Qian et al. (2017) 19.2 16.4 20.0 17.2 18.4 14.2 13.7 11.8 11.5 6.5 12.7 8.5 15.9 12.9
HACNN Li et al. (2018) 12.5 11.1 20.5 19.4 16.7 13.3 9.2 6.2 6.3 2.2 14.8 12.0 13.3 10.7
MLFN Chang et al. (2018) 20.2 18.9 21.1 19.7 17.6 11.1 13.2 12.1 8.3 3.5 13.1 9.1 15.6 12.4
PCB Sun et al. (2018) 23.6 24.2 24.4 25.1 19.9 14.7 20.6 23.6 11.0 6.8 18.6 14.4 19.7 18.1
OSNet Zhou et al. (2019) 19.8 17.3 21.0 19.0 18.7 14.6 12.3 10.9 9.4 5.4 13.0 10.2 15.7 12.9
Multi PFNet Zheng et al. (2021) - - 31.9 29.8 25.5 25.8 - - - - 26.4 23.4 - -
DENet Zheng et al. (2023a) - - 35.4 36.8 33.0 35.4 - - - - 32.4 29.2 - -
TOP-ReID Wang et al. (2024c) 54.4 57.5 64.3 67.6 51.9 54.5 35.3 35.4 26.2 26.0 34.1 31.7 44.4 45.4
@@
Evaluation on Modality-mismatched Scenarios. To evaluate the generalizability of our approach under realistic deployment conditions, we benchmark MDReID across four representative modality-mismatched scenarios on three public datasets. We compare against two strong baselines, TOP-ReID Wang et al. (2024c) and EDITOR Zhang et al. (2024), by reproducing their results using official open-source implementations. As shown in Table 3, EDITOR is designed exclusively for modality-matched settings, which exhibits limited effectiveness when modalities differ across query and gallery. Although TOP-ReID attempts to address missing modalities via reconstruction, its performance remains constrained due to the ill-posed nature of the reconstruction problem. In contrast, our MDReID framework consistently improves the average mAP by 3.4%, 11.8%, and 10.9% on the RGBNT201, RGBNT100, and MSVR310 datasets, respectively. In addition, we train four expert models on the RGBNT201 dataset for four specific scenarios. Results show that our method outperforms these experts by 9.2% on average in mAP and by 8.9% in R‑1. Importantly, our single‑model approach adapts to all four scenarios and offers greater flexibility. In summary, these results highlight the robustness and adaptability of MDReID in handling modality-mismatched ReID, affirming its effectiveness capable of supporting arbitrary modality combinations in both query and gallery inputs.
@@
RGBNT201 EDITOR Zhang et al. (2024) 27.3 27.9 2.80 0.0 4.0 2.3 4.3 4.1 8.5 7.5
TOP-ReID Wang et al. (2024c) 43.0 44.6 14.4 13.3 15.4 14.0 11.9 8.7 18.2 18.0
@@
RGBNT100 EDITOR Zhang et al. (2024) 42.1 59.9 2.6 0.8 2.8 1.5 2.7 1.5 11.9 15.5
TOP-ReID Wang et al. (2024c) 59.0 81.8 21.9 28.5 26.2 34.0 21.7 25.4 26.8 36.1
@@
MSVR310 EDITOR Zhang et al. (2024) 6.4 11.0 2.2 2.5 1.6 0.2 1.7 1.5 2.5 3.4
TOP-ReID Wang et al. (2024c) 18.4 30.5 12.9 19.5 13.7 21.3 12.6 18.8 11.2 17.8